OpenStudy (anonymous):

A ship sails from a harbor for 16 nautical miles south. It then sails in the direction S 55 E for 18 nautical miles. How far is it then from the harbor to the nearest nautical mile? 16 nautical miles 23 nautical miles 30 nautical miles

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

What we have is a triangle with one side 16 and the other side 18, and the angle between them is (180 - 55) = 125. We can use the Law of Cosines to solve for the unknown side, which will be the distance from the boat to the harbor. c^2 = a^2 + b^2 - 2ab*cosC c = sqrt( 16^2 +18^2 - 2*16*18*cos(125) ) c = 30.172, so the answer is 30 nautical miles.

OpenStudy (anonymous):

omg thanks

OpenStudy (anonymous):

Your welcome :)